if f(x)=x² and g(x)=2x-1.find1.gf(x) 2.fg(x) 3.inverse of g(x)?


1
Expert's answer
2022-02-22T13:37:01-0500

1.

"(g\circ f)(x)=2x^2-1"

2.


"(f\circ g)(x)=(2x-1)^2"

3.


"g(x)=2x-1"

Replace "g(x)" with "y"


"y=2x-1"

Replace "y" with "x"


"x=2y-1"

Solve for "y"


"y=\dfrac{1}{2}x+\dfrac{1}{2}"

Replace "y" with "g^{-1}(x)"


"g^{-1}(x)=\dfrac{1}{2}x+\dfrac{1}{2}"